OpenStudy (anonymous):

Solve: square root x+4=-10 square root x=4 square root x+4=10

OpenStudy (kira_yamato):

sqrt(x+4) = -10 or sqrt(x) + 4 = -10?

OpenStudy (anonymous):

sqrt x+4=-10

OpenStudy (anonymous):

Is it : \[\sqrt{x+4} = 10\] or \[\sqrt{x}+4=-10\]? it's very different. sqrt(9+16) = 5 where sqrt(9) + 16 = 19